A teacher asked her class to build a model pyramid with a square base. The base was to have an area of 64 square centimeters and the height of the triangular faces was to be 12 centimeters.What would the surface area of the pyramid be?
Answer by mananth(12270)

A teacher asked her class to build a model pyramid with a square base. The base was to have an area of 64 square centimeters and the height of the triangular faces was to be 12 centimeters.What would the surface area of the pyramid be?
...
area of base = 64cm^2
area of 1 traingle = 1/2 * 8*12
=48 cm^2
4 faces
4*28= 112 cm^2
..
Total area = 112 + 64 = 176 cm^2
...
m.ananth@hotmail.ca